# Optimal and Fair Encouragement Policy Evaluation and Learning

### v3

@@ 3 Problem Setup
Our key assumption beyond standard causal inference assumptions is assumption 2, the conditional exclusion restriction, i.e., that, conditional on observable information $X$, the recommendation has no causal effect on the outcome beyond its effect on the treatment probability. This assumes that all of the covariate information that is informative of downstream outcomes is measured. Although this assumption may not exactly hold in all applications, stating it is also a starting point for sensitivity analysis under violations of it (Kallus and Zhou, 2018; Kallus et al., 2019b).
Assumption 4 is a structural assumption that limits our method to most appropriately reoptimize over small changes to existing algorithmic recommendations. For example, $p_{0\mid 1}(x)$ (disagreement with algorithmic recommendation) could be a baseline algorithmic aversion. Not all settings are appropriate for this assumption. We do not assume microfoundations on how or why human decision-makers deviate from algorithmic recommendations; rather, we take these patterns as given. Again, we can relax this assumption with sensitivity analysis. Assumption 5 is a mild assumption on modeling utility, namely, that it is not defined on joint realizations of potential outcomes.
@@ 4 Method
$R$ is (as-if) randomized and satisfies overlap (Assumption 6).
Then, $R$ can be interpreted as the ITT or prescription, whereas $T$ is the actual realization thereof. 1 models nonadherence situations where decision-makers can target encouragements but not direct receipt of the treatment itself.
@@
$R$ is the output of a predictive model and does not satisfy Assumption 6.
@@
Proposition 2 (Threshold solutions).
@@
We retain the full expression rather than simplifying (as appears in Qiu et al., 2021) since the doubly robust estimation of constraints changes the Lagrangian. For example, for regression adjustment, it is clearer in Proposition 9 how constraints affect the optimal decision rule.
@@ 4.1 Robust estimation with treatment overlap but without recommendation overlap
When the recommendations are, e.g., the high risk/low risk labels from binary classifiers, the overlap assumption may not be satisfied since the algorithmic recommendations are deterministic functions of covariates. However, note that identification in Proposition 1 requires only SUTVA, consistency, and the exclusion restriction.
A naive approach based on parametric extrapolation is to estimate $p_{1\mid 1}(X)$, treatment responsivity, on the observed data and simply use the parametric form to extrapolate to the full dataset. (In Appendix B, we describe the variance reduction that can be possible). On the other hand, parametric extrapolation is generally unsatisfactory because the conclusions will be driven by the model specification rather than by observed data. Nonetheless, it can provide a starting point for robust extrapolation of structurally plausible treatment response probabilities.
@@
We next describe methods for robust extrapolation under structural assumptions about the smoothness of the outcome models. Under violations of overlap, the only unknown quantity is $p_{t\mid r}(X)$ in regions of no overlap in recommendation; however, a plausible assumption is that the underlying function is smooth in covariates. A robust approach obtains worst-case bounds on policy value under all functions compatible with a particular smoothness assumption. On the other hand, we assume that overlap holds with respect to $T$ given covariates $X$, so our finer-grained approach via Assumption 2 yields milder penalties due to robustness since we need robustly extrapolate only the treatment response to recommendations, $p_{t\mid r}(X),$ rather than the outcome model’s $\mu_{t}(X).$ Define the regions of no overlap as the following: Let $\mathcal{X}^{\text{nov}}_{r}=\{x:P(R=r\mid x)=0\}$; in this region, we do not jointly observe all potential values of $(t,r,x)$. In addition, let $\mathcal{X}^{\text{nov}}=\bigcup_{r}\mathcal{X}^{\text{nov}}_{r}$. Correspondingly, define the overlap region as $\mathcal{X}^{\text{ov}}=(\mathcal{X}^{\text{nov}})^{c}.$ We consider uncertainty sets for ambiguous treatment recommendation probabilities. For example, one plausible structural assumption is monotonicity of treatment in recommendation. We define the following uncertainty set:
@@
[…]
@@
We consider the case of continuous-valued outcomes in the example setting of the simple resource parity–constrained program of Equation 2. We first study simple uncertainty sets, such as intervals, to deduce insights about the robust policy, with a more general reformulation in the appendix.
@@ 5 Additional Fairness Constraints and Policy Optimization
The elements of $h_{j}(\pi)$ are average functionals (for example, the average treatment takeup in group $j$). Importantly, the moment function $g_{j}$ depends on $\pi$, while the conditioning event $\mathcal{E}_{j}$ cannot depend on $\pi$. Many important fairness constraints can nonetheless be written in this framework, such as burden/resource parity and parity in true positive rates, but not measures such as calibration whose conditioning event does depend on $\pi$. (See Section C.2 for examples omitted for brevity.)
@@
We use a well-known reduction of policy learning to cost-sensitive classification, described in Section C.2.1 of the appendix. Therefore, the centered regret can be reparametrized via the parameter $\beta$ as: $J(\beta)=J(\operatorname{sgn}(f_{\beta}(\cdot)))=\mathbb{E}[\operatorname{sgn}(f_{\beta}(X))\left\{\psi\right\}].$ We can apply the standard reduction to cost-sensitive classification since $\psi_{i}\operatorname{sgn}(f_{\beta}(X_{i}))=\left|\psi_{i}\right|(1-2\mathbb{I}\left[\operatorname{sgn}(f_{\beta}(X_{i}))\neq\operatorname{sgn}(\psi_{i})\right]$). Then, we can use surrogate losses for the zero-one loss. Although many functional forms for $\ell(\cdot)$ are Fisher consistent, one such choice of $\ell$ is the logistic (cross-entropy) loss $\mathbb{E}[\left|\psi\right|\ell(f_{\beta}(X),\operatorname{sgn}(\psi))],\qquad l(g,s)=2\log(1+\exp(g))-(s+1)$.
@@
2:  Obtain $\hat{Q}^{*}_{1}$ and the index set of binding constraints ^$\hat{\mathcal{I}}_{1}$ by learning nuisances $\eta_{1}$ and running Algorithm 1 on $\mathcal{D}_{1}$ with $\text{REDFAIR}(\mathcal{D}_{1},h,{\mathcal{E}},M,d;\eta_{1}).$
3:  $\hat{\sigma}^{2}_{j}\leftarrow\hat{\operatorname{Var}}(g_{j}(O,\hat{Q}_{1})\mathbb{I}\left[\mathcal{E}_{j}\right]/p_{j}),\;\forall j$ $\hat{d}\leftarrow d+2\sum_{j\in\mathcal{J}}\left|M_{k,j}\right|\hat{\sigma}^{2}_{j}n^{-\alpha}$
4:  Augment additional constraints with $\epsilon_{n}$ policy value and constraint slices relative to $\hat{\pi}_{1}$; define an augmented system (where subscripting by ^$\hat{\mathcal{I}}_{1}$ subindexes the corresponding matrix or vector):
@@
5:  Solve $\min_{Q\in\Delta(\Pi)}\{\hat{V}(Q)\colon\;\;\tilde{M}\tilde{h}(Q)\leq\tilde{d}\}.$ Obtain $\hat{Q}^{*}_{2}$ by running Algorithm 1 on $\mathcal{D}_{2}$ with $\text{REDFAIR}(\mathcal{D},\tilde{g},\tilde{\mathcal{E}},\tilde{M},\tilde{d},\eta_{2})$.
@@
Next, we summarize the optimization algorithm. We play a no-regret (second-order multiplicative weights; Cesa-Bianchi et al., 2007; Steinhardt and Liang, 2014) algorithm (a slight variant of the hedge/exponentiated gradient algorithm; Freund and Schapire, 1997) for the $\lambda$-player while using best-response oracles for the $Q$-player. Full details are in Algorithm 1. Given $\lambda_{t},$ $\operatorname{BEST}_{\beta}\left({\lambda}_{t}\right)$ computes a best response over $Q$; since the worst-case distribution will place all its weight on one classifier, this step can be implemented by a reduction to cost-sensitive/weighted classification (Beygelzimer and Langford, 2009; Zhao et al., 2012), which we describe in further detail below. Computing the best response over $\operatorname{BEST}_{{\lambda}}(\hat{Q}_{t}))$ selects the most violated constraint. Further details are in Section C.2.
@@
These challenges motivate the two-stage procedure, described formally in Algorithm 2 and verbally here. We adapt an out-of-sample regularization scheme developed in (Chernozhukov et al., 2019), which recovers variance-sensitive regret bounds via a small modification to an ERM procedure (and, by extension, policy learning). We split the data into two subsets, $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, and first learn nuisance estimators $\hat{\eta}_{1}$ from $\mathcal{D}_{1}$ (possibly with further sample splitting) for use in our policy value and constraint estimates. We run Algorithm 1 ($\operatorname{REDFAIR}(\mathcal{D}_{1},h,{\mathcal{E}},M,d;\hat{\eta}_{1})$) on data from $\mathcal{D}_{1}$ to estimate the optimal policy distribution $\hat{Q}_{1}$ and the constraint variances at $\hat{Q}_{1}$. We identify the first-stage binding constraints via the index set $\hat{I}_{1}$. Next, we augment the constraint matrix with additional constraints that require feasible policies for the second-stage policy distribution to achieve $\epsilon_{n}$ close policy value and constraint moment values relative to $\hat{Q}_{1}$. Since errors concentrate quickly, this can be viewed as variance regularization. In addition, we set the constraint slacks $\hat{d}\leftarrow d+2\sum_{j\in\mathcal{J}}\left|M_{k,j}\right|\hat{\sigma}^{2}_{j}n^{-\alpha}$ in the second stage using the estimated variance constants from $\hat{Q}_{1}$. This results in tighter control of the fairness constraints. The second stage solves for an approximate saddle point of the augmented system, with objective function and constraints evaluated on $\mathcal{D}_{2}$ and returns $\hat{Q}_{2}.$
[…]

### v4

@@ 2 Related Work
Algorithmic advice in operations
A large and rapidly growing literature in operations studies algorithmic advice, at times from a behavioral focus (but not specialized to causal effects). (Ge et al., 2025) studies classification conditional parity under no-information on compliance, and proposes compliance-robust policies sandwiched in between the range of human decisions (under no structure as to how human responders behave). (Grand-Clément and Pauphilet, 2024) also shares a focus on leveraging adherence to optimize advice rather than decisions - however they adopt a dynamic model with updating. We focus on implications for optimizing causal encouragements as well. (McLaughlin and Spiess, 2022; Gillis et al., 2021) study the fairness of machine-assisted human decisions and develop a principal-agent model studying how algorithmic advice might shift beliefs in risk and preferences. In contrast, we are not micro-founded as to why disparities in compliance arise, which is an important direction of future work. On the non-algorithmic advice, but encouragement side, (Freund and Hssaine, 2025) consider dynamic resource allocation under fairness constraints, when deciding monetary incentives for retention under stochastic participation. Other models of stochastic resource usage and demand are also relevant, though potentially more specialized than our non-adherence setting.
@@ 3 Problem Setup
Our key assumption beyond standard causal inference assumptions is assumption˜2, the conditional exclusion restriction, i.e., that, conditional on observable information $X$, the recommendation has no causal effect on the outcome beyond its effect on the treatment probability. This assumes that all of the covariate information that is informative of downstream outcomes is measured. Although this assumption may not exactly hold in all applications, stating it is also a starting point for sensitivity analysis under violations of it (Kallus and Zhou, 2018; Kallus et al., 2019b).
Assumption˜4 is a structural assumption that limits our method to most appropriately reoptimize over small changes to existing algorithmic recommendations. For example, $p_{0\mid 1}(x)$ (disagreement with algorithmic recommendation) could be a baseline algorithmic aversion. Not all settings are appropriate for this assumption. We do not assume microfoundations on how or why human decision-makers deviate from algorithmic recommendations; rather, we take these patterns as given. Again, we can relax this assumption with sensitivity analysis. Assumption˜5 is a mild assumption on modeling utility, namely, that it is not defined on joint realizations of potential outcomes.
@@
$R$ is (as-if) randomized and satisfies overlap (Assumption˜6).
Then, $R$ can be interpreted as the ITT or prescription, whereas $T$ is the actual realization thereof. ˜1 models nonadherence situations where decision-makers can target encouragements but not direct receipt of the treatment itself.
@@
$R$ is the output of a predictive model and does not satisfy Assumption˜6.
@@
4 Analysis: What drives utility and budget unfairness under naive ITT targeting?
In order to motivate our later methodological work on estimating and optimizing encouragement policies under fairness constraints, we first answer the question: why is naive ITT targeting insufficient to understand the key drivers of unfairness? We focus on the class of threshold policies under global budget constraints. (Targeting encouragements is only needed when there is a budget constraint, as it is rare for encouragements to have negative treatment effects). We highlight that explicitly disaggregating effects on enrollment into treatment enables isolating the source of disparities into two substantively distinct mechanism: outreach vs. interventional effectiveness. We illustrate with a preview of our case study on SNAP outreach how unfairness surfaces in real-world scenarios. We develop analytical fairness decompositions that highlight key diagnostic metrics that lower-bound disparities in policy value. These metrics can be estimated from data — we highlight these in the case of the SNAP case study.
We begin by considering naive targeting on the intention-to-treat effect, i.e. viewing the encouragement as a treatment itself. Define the encouragement score:
$s(X,A)\coloneqq\tau(X,A)\,(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A)).$
Proposition 1 (Quantile-threshold optimality).
Assume the distribution of $s(X,A)$ is continuous (no atoms). Consider optimizing the ITT effect of encouragements under a global encouragement budget:
$\max_{\pi}\{\mathbb{E}[Y(\pi)]\colon\mathbb{E}[\pi(X,A)]\leq b\}$ (encouragement budget)
Let $q_{s}:[0,1]\to\mathbb{R}$ denote the (right–continuous) quantile function of $s$, $q_{s}(t)\;\coloneqq\;\inf\{z\in\mathbb{R}:\;\mathbb{P}\big(s(X,A)\leq z\big)\geq t\}$ Then the optimal policy is a threshold policy at the top-$b$ quantile of the encouragement score:
$\pi_{b}(x,a)\;=\;\mathbf{1}\!\left\{\,s(x,a)\geq\max(q_{s}(1-b),0)\,\right\}.$
Unfairness arises from differences in marginal levels, and joint dependence between treatment effect and nudgeability.
Fix a baseline policy $\pi_{0}$. The (group-conditional) improvement of $\pi$ over $\pi_{0}$ is
${Imp}_{a}(\pi;\pi_{0}):=\mathbb{E}\!\left[Y(\pi)-Y(\pi_{0})\mid A=a\right]=\mathbb{E}\!\left[\tau_{a}\,c_{a}\,(\pi(X,a)-\pi_{0}(X,a))\mid A=a\right]$
Improvement and disparity at a single threshold.
Choose $\pi_{0}=0$, i.e. encourage no-one. For example, this could be the status quo without any encouragements. Then
${Imp}_{a}(\pi^{*};0)=\mathbb{E}\!\left[s_{a}\,\mathbb{I}\{s_{a}\geq t^{*}\}\mid A=a\right],$
and analogously for $A=b$. Let the induced within-group budget be $\beta_{a}^{t^{*}}:=P(s_{a}\geq t^{*}\mid A=a)$.
The improvement at a threshold policy is driven by marginal levels of thresholded $\tau,c$ as well as their conditional covariance:
${Imp}_{a}(\pi^{*};0)={\mathbb{E}[\tau_{a}\mid s_{a}\geq t^{*}]\ \mathbb{E}[c_{a}\mid s_{a}\geq t^{*}]}+\operatorname{Cov}\!\Big(\tau_{a}-\mathbb{E}[\tau_{a}\mid s_{a}\geq t^{\star}],\ c_{a}-\mathbb{E}[c_{a}\mid s_{a}\geq t^{\star}]\,\Big|\ s_{a}\geq t^{\star}\Big)$
The conditional covariance term reflects the contribution of joint dependence of the treatment effect and compliance effects. It is zero (within the encouraged stratum) iff the treatment effect and compliance are uncorrelated, and positive when higher-benefit units also tend to comply more (as a “win-win" situation). Targeting encouragements is particularly beneficial when the people who benefit most are also the ones where encouragement nudges into treatment. Misalignment between effectiveness and nudgeability, surfacing as low or even negative dependence, suggest that encouragements need to be redesigned to achieve better targeting.
The disparity in improvement at $t^{*}$ is
$\Delta^{\mathrm{imp}}(t^{*}):={Imp}_{a}(t^{*})-{Imp}_{b}(t^{*})=\beta_{a}^{t^{*}}\mathbb{E}\!\left[s_{a}\,\mathbb{I}\{s_{a}\geq t^{*}\}\mid A=a\right]-\beta_{b}^{t^{*}}\mathbb{E}\!\left[s_{b}\,\mathbb{I}\{s_{b}\geq t^{*}\}\mid A=b\right]$
In the general case, the presence of disparity will depend on the budget (and therefore threshold): where in each group’s score distribution does the threshold fall, and what is the dependence structure of heterogeneity and nudgeability for each group. Differing strengths of dependence across groups can also result in disparities in group-specific utilities. The fundamental object that describes such dependence is the copula, the bivariate joint distribution between each group’s distribution of heterogeneous treatment effect and compliance effect functions. A general scale-free measure of dependence is the Spearman’s rank correlation, which measures the correlation of the ranks of two random variables and hence indicates whether a monotonic (but not necessarily linear) relationship exists between them. Under stronger structure on the copula, one can leverage disparities in metrics like Spearman’s rank correlation to conclude that disparities would persist uniformly over all potential thresholds. The required condition is a concordance-ordered copula family assumption, which posits that each group’s joint distribution of $(\tau,c)$ is indexed by a single dependence parameter $\theta$, such that larger $\theta$ values indicate stronger positive dependence—greater alignment between individuals who benefit most and those most likely to comply. Examples include the Gaussian or Archimedean (Gumbel, Clayton) copulas.
Theorem 3 (Spearman lower bound for disparity at the pooled cut).
Assume that for each group the copula belongs to a concordance-ordered family, indexed by a parameter $\theta,$ $\{C_{\theta}:\theta\in\Theta\}$, and that Spearman’s rank correlation $\rho_{\mathrm{Sp}}(\tau_{a},c_{a})$ (resp. $\rho_{\mathrm{Sp}}(\tau_{b},c_{b})$) is strictly increasing in $\theta$. Then there exist strictly positive constants $\kappa(t^{*})$, depending on the groupwise score distributions and copulas, so that
$\Delta^{\mathrm{imp}}(t^{*})\ \geq\ \kappa_{*}(t^{*})\ (\rho_{\mathrm{Sp}}^{\,a}-\rho_{\mathrm{Sp}}^{\,b})\ +(\mathbb{E}[\tau_{a}\mid s_{a}\geq t^{*}]\mathbb{E}[c_{a}\mid s_{a}\geq t^{*}]-\mathbb{E}[\tau_{b}\mid s_{b}\geq t^{*}]\mathbb{E}[c_{b}\mid s_{b}\geq t^{*}])\ .$
In short: under concordance-ordered copulas, if one group has consistently stronger rank-alignment between treatment effect and compliance, then regardless of the budget threshold, disparities in policy value improvement persist. Such disparities could point to the need for more intensive outreach or differently designed encouragements.
4.1 Case study: Text message reminders for SNAP recertification
Background
Each year, households receiving SNAP (food assistance) must complete a short recertification interview to verify continued eligibility. Missing this interview results in termination of benefits, even for households that remain eligible.
We study a pilot program in San Francisco in which clients could opt in to receive a text-message reminder about the recertification deadline. The reminder encouraged participants to complete the required interview—our effective treatment—so that their benefits would continue. Since many who fail to recertify are in fact eligible, increasing interview attendance often improves benefit continuity and household welfare.
Using data from Homonoff and Somerville (2021), we reinterpret the reminder as an encouragement $R$ affecting the take-up of treatment $T$ (interview completion), which in turn influences the outcome $Y$ (next-year SNAP benefits). We focus on heterogeneity in both the reminder’s compliance effect and the treatment effect of interview completion, with particular attention to racial disparities in their alignment.
Text reminders are a low-cost operational lever that can reduce administrative burdens, but limited budgets may constrain how widely they are deployed. Our analysis evaluates how to target such reminders to maximize overall benefit take-up while considering fairness in access and impact across groups.
Descriptives: heterogeneity vs nudgeability.
In Figure˜1, we compare the distributions within each race (binarized to white $A=0$ or nonwhite $A=1$) of the heterogeneous treatment effect $\mathbb{E}[Y(1)-Y(0)\mid X],$ the heterogeneous compliance effect $p_{1\mid 1}(X)-p_{1\mid 0}(X),$ and the heterogeneous encouragement effect, which is the product of the previous two: $(\mathbb{E}[Y(1)-Y(0)\mid X])(p_{1\mid 1}(X)-p_{1\mid 0}(X))$. In a setting with fixed costs of recommendation and treatment, the default approach of viewing the recommendation as the treatment, i.e., an optimal unconstrained encouragement regime, sorts by the latter score.
In this setting, we find that treatment efficacy is broadly correlated with takeup: Higher treatment effects are rank correlated with higher compliance effects. However, the strength of this relationship, quantified via the Spearman’s $\rho_{SP}$ rank correlation coefficient, varies across groups, leading to inequality. Among white beneficiaries only, the rank correlation coefficient between heterogeneous treatment and compliance effects is 0.68, while it is 0.45 for nonwhite beneficiaries. While both correlations are statistically significantly positive, the relationship is weaker for nonwhite than for white beneficiaries, in particular with regard to those most likely to take up.
A central concern here is that worse-off groups might have more to gain by accessing the treatment (i.e., have heterogeneous treatment effects of greater magnitude) but may respond less to nudges/recommendations into treatment (because of behavioral frictions, time/resource constraints, higher need, etc.). In this setting, decreasing marginal utilities of money establish the former, while the process of filling out the burdensome benefits application or making it to a recertification interview under time, childcare, transportation, or resource constraints speaks to the relative difficulty posed by the latter. The density plots in Figure˜1 underscore this phenomenon. The first plot shows the treatment effect heterogeneity: The distribution is wider for nonwhite than for white beneficiaries; i.e., it has more mass at higher magnitudes of $\tau$ and less at lower magnitudes of $\tau$. However, we see the opposite effect on compliance: While the results are relatively weak, they are bimodal, with a substantially greater probability density at higher compliance treatment effects for white than for nonwhite beneficiaries. One reason for this could be the language barrier for non-English-speaking applicants.
Our final analysis considers equity in allocation efficiency, specifically the lift or improvements to decisions around who receives recommendations. In Figure˜3, we evaluate the lift in different performance metrics conditional on protected group and conditional on the group’s being encouraged under a resource-constrained policy—the y-axes show $\mathbb{E}[c(Y(\pi)))\mid A=a,\pi^{*}(X)=1],\mathbb{E}[(Y(1)-Y(0))\mid A=a,\pi^{*}(X)=1],\mathbb{E}[(T(1)-T(0))\mid A=a,\pi^{*}(X)=1]$. The x-axis ranges over thresholds on the conditional encouragement effect, $(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\times\tau(X,A)$. Thus, the x-axis explores the range of the population covered under the budget allocation, with the budget averaged over the full population on the left-hand side and encouragement restricted to few individuals on the right-hand side. This represents the range of resource-constrained rules under a naive ITT approach. We do find, for large ranges of moderate budgets, that the naive efficient allocation results in disparities in average outcomes for white and nonwhite beneficiaries. The next two figures on the right disaggregate the individual encouragement effect into the heterogeneous causal effect $\tau$ of interview attendance and the compliance effect $(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))$ of text message reminders on interview attendance. Consistent with the earlier results on the strong positive rank correlation of $\tau$ and $(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))$, both plots are nearly nondecreasing as we threshold based on increasing individual encouragement scores $(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\times\tau$. However, as we also saw earlier, this correlation is weaker for nonwhite beneficiaries, and we can observe a small region of nonmonotonic behavior.
Comparing these two plots, we see that the disparities in naive ITT-based allocation are in fact driven almost entirely by disparities in access, which would persist under a large range of budgets. While leveraging disparities in Spearman’s correlation to diagnose disparities in resulting policy values is scale-free, Figure˜3 exactly highlights the magnitude of the conditional covariance gap.
Disaggregating our analysis by looking separately at treatment vs. compliance effects and at covariate-heterogeneous effects allows us to quantify access gaps that would persist even under efficient allocations. Isolating these sources supports firms in what to do next to reduce disparities, even if it may not be the fairness-constrained methods we next study in generality. Considering treatment effects and compliance effects separately is crucial since they imply focusing on different strategies moving forward, or because one may be outside of the firm’s control. For example, in the SNAP case study, only outreach and administrative (like how to schedule interviews) mechanisms have room for freedom for local organizations.
5 Method
@@
Proposition 3 (Threshold solutions under resource constraints).
@@
We retain the full expression rather than simplifying (as appears in Qiu et al., 2021) since the doubly robust estimation of constraints changes the Lagrangian. For example, for regression adjustment, it is clearer in Proposition˜10 how constraints affect the optimal decision rule.
@@ 5.1 Robust estimation with treatment overlap but without recommendation overlap
When the recommendations are, e.g., the high risk/low risk labels from binary classifiers, the overlap assumption may not be satisfied since the algorithmic recommendations are deterministic functions of covariates. However, note that identification in Proposition˜2 requires only SUTVA, consistency, and the exclusion restriction.
A naive approach based on parametric extrapolation is to estimate $p_{1\mid 1}(X)$, treatment responsivity, on the observed data and simply use the parametric form to extrapolate to the full dataset. (In Appendix˜B, we describe the variance reduction that can be possible). On the other hand, parametric extrapolation is generally unsatisfactory because the conclusions will be driven by the model specification rather than by observed data. Nonetheless, it can provide a starting point for robust extrapolation of structurally plausible treatment response probabilities.
@@
We next describe methods for robust extrapolation under structural assumptions about the smoothness of the outcome models. Under violations of overlap, the only unknown quantity is $p_{t\mid r}(X)$ in regions of no overlap in recommendation; however, a plausible assumption is that the underlying function is smooth in covariates. A robust approach obtains worst-case bounds on policy value under all functions compatible with a particular smoothness assumption. On the other hand, we assume that overlap holds with respect to $T$ given covariates $X$, so our finer-grained approach via Assumption˜2 yields milder penalties due to robustness since we need robustly extrapolate only the treatment response to recommendations, $p_{t\mid r}(X),$ rather than the outcome model’s $\mu_{t}(X).$ Define the regions of no overlap as the following: Let $\mathcal{X}^{\text{nov}}_{r}=\{x:P(R=r\mid x)=0\}$; in this region, we do not jointly observe all potential values of $(t,r,x)$. In addition, let $\mathcal{X}^{\text{nov}}=\bigcup_{r}\mathcal{X}^{\text{nov}}_{r}$. Correspondingly, define the overlap region as $\mathcal{X}^{\text{ov}}=(\mathcal{X}^{\text{nov}})^{c}.$ We consider uncertainty sets for ambiguous treatment recommendation probabilities. For example, one plausible structural assumption is monotonicity of treatment in recommendation. We define the following uncertainty set:
@@
$\text{ where }\overline{V}_{nov}(\pi)$ $\textstyle\coloneqq\underset{q_{tr}(X)\in\mathcal{U}}{\max}\left\{\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mu_{t}(X)q_{tr}(X)\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}]]\right\}$
@@
We consider the case of continuous-valued outcomes in the example setting of the simple resource parity–constrained program of LABEL:eqn-resource-parity-constrained. We first study simple uncertainty sets, such as intervals, to deduce insights about the robust policy, with a more general reformulation in the appendix.
@@ 6 Additional Fairness Constraints and Policy Optimization
The elements of $h_{j}(\pi)$ are average functionals (for example, the average treatment takeup in group $j$). Importantly, the moment function $g_{j}$ depends on $\pi$, while the conditioning event $\mathcal{E}_{j}$ cannot depend on $\pi$. Many important fairness constraints can nonetheless be written in this framework, such as burden/resource parity and parity in true positive rates, but not measures such as calibration whose conditioning event does depend on $\pi$. (See Section˜C.2 for examples omitted for brevity.)
@@
We use a well-known reduction of policy learning to cost-sensitive classification, described in Section˜C.2.1 of the appendix. Therefore, the centered regret can be reparametrized via the parameter $\beta$ as: $J(\beta)=J(\operatorname{sgn}(f_{\beta}(\cdot)))=\mathbb{E}[\operatorname{sgn}(f_{\beta}(X))\left\{\psi\right\}].$ We can apply the standard reduction to cost-sensitive classification since $\psi_{i}\operatorname{sgn}(f_{\beta}(X_{i}))=\left|\psi_{i}\right|(1-2\mathbb{I}\left[\operatorname{sgn}(f_{\beta}(X_{i}))\neq\operatorname{sgn}(\psi_{i})\right]$). Then, we can use surrogate losses for the zero-one loss. Although many functional forms for $\ell(\cdot)$ are Fisher consistent, one such choice of $\ell$ is the logistic (cross-entropy) loss $\mathbb{E}[\left|\psi\right|\ell(f_{\beta}(X),\operatorname{sgn}(\psi))],\qquad l(g,s)=2\log(1+\exp(g))-(s+1)$.
@@
2: Obtain $\hat{Q}^{*}_{1}$ and the index set of binding constraints $\hat{\mathcal{I}}_{1}$ by learning nuisances $\eta_{1}$ and running Algorithm˜1 on $\mathcal{D}_{1}$ with $\text{REDFAIR}(\mathcal{D}_{1},h,{\mathcal{E}},M,d;\eta_{1}).$
[…]
4: Augment additional constraints with $\epsilon_{n}$ policy value and constraint slices relative to $\hat{\pi}_{1}$; define an augmented system (where subscripting by $\hat{\mathcal{I}}_{1}$ subindexes the corresponding matrix or vector):
@@
5: Solve $\min_{Q\in\Delta(\Pi)}\{\hat{V}(Q)\colon\;\;\tilde{M}\tilde{h}(Q)\leq\tilde{d}\}$. Obtain $\hat{Q}^{*}_{2}$ by running Algorithm˜1 on $\mathcal{D}_{2}$ with $\text{REDFAIR}(\mathcal{D},\tilde{g},\tilde{\mathcal{E}},\tilde{M},\tilde{d},\eta_{2})$.
@@
Next, we summarize the optimization algorithm. We play a no-regret (second-order multiplicative weights; Cesa-Bianchi et al., 2007; Steinhardt and Liang, 2014) algorithm (a slight variant of the hedge/exponentiated gradient algorithm; Freund and Schapire, 1997) for the $\lambda$-player while using best-response oracles for the $Q$-player. Full details are in Algorithm˜1. Given $\lambda_{t},$ $\operatorname{BEST}_{\beta}\left({\lambda}_{t}\right)$ computes a best response over $Q$; since the worst-case distribution will place all its weight on one classifier, this step can be implemented by a reduction to cost-sensitive/weighted classification (Beygelzimer and Langford, 2009; Zhao et al., 2012), which we describe in further detail below. Computing the best response over $\operatorname{BEST}_{{\lambda}}(\hat{Q}_{t}))$ selects the most violated constraint. Further details are in Section˜C.2.
@@
These challenges motivate the two-stage procedure, described formally in Algorithm˜2 and verbally here. We adapt an out-of-sample regularization scheme developed in (Chernozhukov et al., 2019), which recovers variance-sensitive regret bounds via a small modification to an ERM procedure (and, by extension, policy learning). We split the data into two subsets, $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, and first learn nuisance estimators $\hat{\eta}_{1}$ from $\mathcal{D}_{1}$ (possibly with further sample splitting) for use in our policy value and constraint estimates. We run Algorithm˜1 ($\operatorname{REDFAIR}(\mathcal{D}_{1},h,{\mathcal{E}},M,d;\hat{\eta}_{1})$) on data from $\mathcal{D}_{1}$ to estimate the optimal policy distribution $\hat{Q}_{1}$ and the constraint variances at $\hat{Q}_{1}$. We identify the first-stage binding constraints via the index set $\hat{I}_{1}$. Next, we augment the constraint matrix with additional constraints that require feasible policies for the second-stage policy distribution to achieve $\epsilon_{n}$ close policy value and constraint moment values relative to $\hat{Q}_{1}$. Since errors concentrate quickly, this can be viewed as variance regularization. In addition, we set the constraint slacks $\hat{d}\leftarrow d+2\sum_{j\in\mathcal{J}}\left|M_{k,j}\right|\hat{\sigma}^{2}_{j}n^{-\alpha}$ in the second stage using the estimated variance constants from $\hat{Q}_{1}$. This results in tighter control of the fairness constraints. The second stage solves for an approximate saddle point of the augmented system, with objective function and constraints evaluated on $\mathcal{D}_{2}$ and returns $\hat{Q}_{2}.$
Next, we provide a generalization bound on the out-of-sample performance of the policy returned by the two-stage procedure. Importantly, because of our two-stage procedure, the regret of the policy depends on the worst-case variance of near-optimal policies (rather than all policies). Define the function classes $\mathcal{F}_{\Pi}=\{v_{DR}(\cdot,\pi;\eta)\colon\pi\in\Pi,\eta\in\mathcal{F}_{\eta}\}$, $\mathcal{F}_{j}=\{g_{j}(\cdot,\pi;\eta)\colon\pi\in\Pi,\eta\in\mathcal{F}_{\eta}\}$ and the empirical entropy integral $\kappa(r,\mathcal{F})={\inf_{\alpha\geq 0}\{4\alpha+10\int_{\alpha}^{r}\sqrt{\frac{\mathcal{H}_{2}(\epsilon,\mathcal{F},n)}{n}}d\epsilon\}}$, where $H_{2}(\epsilon,\mathcal{F},n)$ is the $L_{2}$ empirical entropy, i.e., log of the $\lVert\cdot\rVert_{2}$ $\epsilon$-covering number. We make a mild assumption of a learnable function class (bounded entropy integral) (Van Der Vaart et al., 1996), which is satisfied by many standard function classes such as linear models, polynomials, kernel regression, and neural networks (Wainwright, 2019).